# When Swin Transformer meets KANs: An Improved Transformer Architecture for Medical Image Segmentation

### v1

@@ 1 Introduction
Medical image segmentation is essential for delineating anatomical structures and pathological regions, supporting diagnosis, quantitative analysis, and treatment planning. Achieving high segmentation accuracy is crucial as even small errors could affect clinical decisions. However, medical image datasets are often constrained by expensive expert annotations and privacy restrictions, making data efficiency a key challenge. Convolutional Neural Network (CNN)–based models such as U-Net [unet] and its variants [isensee2021nnuNet, valanarasu2022unext, milletari2016vnet] have been widely adopted, exploiting locality and translation invariance. But, their limited receptive fields restrict long-range dependency modeling, which is vital for segmenting objects of varying shapes and sizes. Vision Transformers (ViT) [vit] and ViT-based models [hatamizadeh2022unetr, chen2021transunet] addressed this with self-attention [transformer] to capture global relationships and improve context modeling. Still, the vanilla ViT scales poorly for dense prediction (as tokens grow quadratically with resolution) and limits its practicality for high-resolution images. The Swin Transformer (SwinT) [liu2021swin] and SwinT-based medical image segmentation models [cao2022swinUnet, hatamizadeh2021swinUnetr, tangSelfSupervisedPreTrainingSwin2022, he2023swinunetrV2] mitigated this issue through hierarchical shifted windows to reduce computation costs to near-linear while preserving cross-window context. By extracting multi-scale features and balancing local and global contexts, SwinT-based models have proven highly effective for medical image dense prediction tasks such as segmentation.
@@
In ViT and SwinT, Multi-Layer Perceptrons (MLPs) serve as the default Feed-Forward Networks (FFNs) after self-attention blocks to introduce non-linear transformations and channel mixing. However, their reliance on fixed activations such as GELU or ReLU limits their ability to capture complex functional relationships in medical images, and most of their variants still model patterns largely in a linear and opaque manner. Kolmogorov–Arnold Networks (KANs) [liu2024kan, liu2024kan2] address these limitations by replacing static activations with learnable functional expansions, offering greater expressiveness and interpretability. By parameterizing the base functions along each dimension, KANs bridge the gap between a network’s physical attributes and empirical performance. In computer vision, they enhance representation learning by capturing richer non-linear dependencies [bodner2024convolutionalCKAN], and in medical image analysis, they provide a path toward improved data efficiency and finer anatomical detail modeling. Recent work, such as U-KAN [li2025ukan], has shown the potential of incorporating KAN layers into U-Net type architectures, enhancing expressiveness and data efficiency for medical image segmentation. Although effective, UKAN’s reliance on convolutions limits the modeling of long-range dependencies, while its design of using tokenized-KANs only at the bottleneck stage limits their expressive power.
@@
In this work, we explore integrating KANs into Transformer encoders for medical image segmentation, building a more accurate and data-efficient architecture. This design combines the complementary strengths of both components: self-attention in Transformers to capture long-range context essential for complex anatomical structures, and KANs to enrich the feed-forward pathways with flexible and interpretable non-linear approximations for better representation capacity and data efficiency. Transformers with KANs have been investigated in recent work [wang2025swinkan, rifa2025swinKAT] using hybrid MLP–KAN or KAN-convolution designs, while TransUKAN [wu2024transukan] employs CNNs with a KAN bottleneck. In contrast, we integrate KANs directly as feed-forward layers in a standalone SwinT encoder, making ours the first pure Transformer-based encoder with KANs for medical image segmentation.
@@
It is challenging to scale KANs in deep architectures, as vanilla spline-based KANs are computationally inefficient, memory-intensive, and difficult to train [liu2024kan]. To address this, we adopt rational base functions and the Group Rational KAN (GR-KAN) design from KAT [yang2024kolmogorovKAT] to build a more robust Swin Transformer–based architecture. Our proposed model, UKAST (U-shaped segmentation architecture combining KANs with Swin Transformers), mitigates the computation and training inefficiencies of vanilla KANs, yielding a data-efficient segmentation framework with lower FLOPs and comparable parameters as to state-of-the-art (SOTA) SwinUNETR. Our UKAST achieves SOTA performance across diverse modalities of medical image segmentation benchmarks, consistently surpassing both CNN- and Transformer-based baselines. It delivers improved segmentation accuracy in two 2D and even greater improvements in two 3D datasets with approximately the same computational costs in both FLOPs and model parameters. We demonstrate that UKAST’s data efficiency outperforms that of ViTs and SwinTs across various data-scarce settings in both 2D and 3D tasks.
@@ 2.1 UKAST
Given an input 2D medical image of size $H\times W$ with $C$ number of channels, $\mathcal{X}\in\mathbb{R}^{C\times H\times W}$ (for 3D volumes, we operate on their 2D slices), we divide $\mathcal{X}$ into non-overlapping patches of size $H^{\prime}\times W^{\prime}$ each. Every patch is projected into a $C^{\prime}$-dimensional embedding space via a linear layer, forming a sequence of input tokens. The input tokens are partitioned into local windows of size $M\times M$ each in the UKAST encoder, which comprises hierarchical Swin Transformer blocks with Group-KAN feed-forward layers using rational base functions. Within each window, self-attention (W-MSA) is computed to capture local dependencies, where residual convolutions (RC) [he2023swinunetrV2] are also added. To enable cross-window interactions, the subsequent layer applies shifted window attention (SW-MSA), balancing local modeling with global context aggregation. This hierarchical design progressively extracts multi-scale features, encoding both fine-grained details and broader structural information, as follows.
@@ 2.2 Rational Base Functions
We use rational functions as the base functions for KAN layers following KAT [yang2024kolmogorovKAT], rather than B-splines. Specifically, each edge function $\phi(x)$ is parameterized as a rational over polynomials $P(x),Q(x)$ of orders $m$ and $n$ respectively, using the Safe Padé Activation Unit (PAU) [molina2019PAU], as:
@@ 2.3 Group Rational KANs
To enhance the feed-forward pathways, we replace the conventional MLP blocks (used in SwinUNETR) with Group Rational KANs (GR-KANs) following the design of KAT [yang2024kolmogorovKAT]. Instead of using fixed activations, KAN layers learn rational function expansions, enabling expressive non-linear mappings while retaining GPU efficiency. Group-KAN reduces parameter cost by sharing base functions across neurons. Note that a key limitation of the original KAN is that it requires a unique activation function for each input-output pair, leading to a parameter count of $d_{in}\times d_{out}$ and high computational overhead. To overcome this, GR-KAN divides the input channels $d_{in}$ into $g$ groups, sharing rational function parameters within each group. This reduces the number of unique functions needed from $d_{in}\times d_{out}$ to only $g$, attaining significant savings in both parameters and FLOPs. Formally, given an input vector $\mathbf{x}$, GR-KAN applies
@@ 3 Experimental Setup
Datasets: We evaluate our new method on 4 medical image segmentation datasets spanning different modalities: polyp tissue endoscopy (Kvasir-SEG [jha2019kvasir]), skin lesion dermoscopy (ISIC 2017 [codella2018isic]), abdominal CT (BCV 2016 [landman2015bcv]), and cardiac MRI (MMWHS 2017 [zhuang2019evaluationMMWHS]). The training sets were randomly sampled and contain 700 images, 2000 images, 24 volumes (1,982 2D slices), and 16 volumes (3,834 2D slices), respectively. For data efficacy experiments, we use 10%, 25%, 50%, and 100% of the training data. The corresponding test sets include randomly selected 300 images, 600 images, 6 volumes (603 2D slices), and 4 volumes (973 2D slices).
Baselines: We evaluate three CNN-based models (U-Net [unet], U-Net v2 [peng2025unetv2], UNeXt [valanarasu2022unext]), two Transformer-based models (UNETR [hatamizadeh2022unetr], SwinUNETR [tangSelfSupervisedPreTrainingSwin2022]), and one KAN-based model (UKAN [li2025ukan]). We also construct a baseline, UKAT, by replacing the ViT [vit] encoder in UNETR with the KAT encoder [yang2024kolmogorovKAT]. All models are implemented in their smallest or comparable configurations to ensure fairness. For SwinUNETR and our proposed UKAST, we additionally incorporate residual convolution blocks in the encoder stages as used in [he2023swinunetrV2].
Implementation Details: The implementations utilize PyTorch [paszke2019pytorch] and MONAI [cardoso2022monai]. The models were trained from scratch using the AdamW optimizer [loshchilov2017decoupledAdamW] (initial learning rate = 0.0002, weight decay = 0.001) with a cosine annealing learning-rate schedule. The training used a combined Dice and cross-entropy loss on a single NVIDIA A10 GPU for 400 epochs with a batch size of 24. During training, $320\times 320$ random crops were extracted and augmented with horizontal and vertical flips, random $90^{\circ}$ rotations, and Gaussian noise, with masks transformed identically. At inference, predictions were generated using overlapping patches with 50% overlap.

### v2

# Index Terms—  Medical Image Segmentation, Vision Transformers, CNN, Kolmogorov–Arnold Networks

@@ 1 Introduction
Medical image segmentation is essential for delineating anatomical structures and pathological regions, supporting diagnosis, quantitative analysis, and treatment planning. Achieving high segmentation accuracy is crucial as even small errors could affect clinical decisions. However, medical image datasets are often constrained by expensive expert annotations and privacy restrictions, making data efficiency a key challenge. Convolutional Neural Network (CNN)–based models such as U-Net [20] and its variants [10, 24, 16] have been widely adopted, exploiting locality and translation invariance. But, their limited receptive fields restrict long-range dependency modeling, which is vital for segmenting objects of varying shapes and sizes. Vision Transformers (ViT) [6] and ViT-based models [8, 4, 21] addressed this with self-attention [25] to capture global relationships and improve context modeling. Still, the vanilla ViT scales poorly for dense prediction (as tokens grow quadratically with resolution) and limits its practicality for high-resolution images. The Swin Transformer (SwinT) [14] and SwinT-based image segmentation models [2, 7, 23, 9, 22] mitigated this issue through hierarchical shifted windows to reduce computation costs to near-linear while preserving cross-window context. By extracting multi-scale features and balancing local and global context, SwinT-based models have proven highly effective for medical image dense prediction tasks such as segmentation.
@@
In ViT and SwinT, Multi-Layer Perceptrons (MLPs) serve as the default Feed-Forward Networks (FFNs) to introduce non-linear transformations and channel mixing. However, their reliance on fixed activations such as GELU or ReLU limits their ability to capture complex functional relationships in medical images, and most of their variants still model patterns largely in a linear and opaque manner. Kolmogorov–Arnold Networks (KANs) [15] address these limitations by replacing static activations with learnable functional expansions, offering greater expressiveness and interpretability. By parameterizing the base functions along each dimension, KANs bridge the gap between a network’s physical attributes and empirical performance. In computer vision, they enhance representation learning by capturing richer non-linear dependencies [1], and in medical image analysis, they provide a path toward improved data efficiency and finer anatomical detail modeling. Recent work, such as U-KAN [13], has shown the potential of incorporating KAN layers into U-Net type architectures, enhancing expressiveness and data efficiency for medical image segmentation. Although effective, UKAN’s reliance on convolutions limits the modeling of long-range dependencies, while its design of using tokenized-KANs only at the bottleneck stage limits their expressive power.
@@
In this work, we explore integrating KANs into Transformer encoders for medical image segmentation, building a more accurate and data-efficient architecture. This design combines the complementary strengths of both components: self-attention in Transformers to capture long-range context essential for complex anatomical structures, and KANs to enrich the feed-forward pathways with flexible and interpretable non-linear approximations for better representation capacity and data efficiency. Transformers with KANs have been investigated in recent work [26, 19] using hybrid MLP–KAN or KAN-convolution designs, while TransUKAN [27] employs CNNs with a KAN bottleneck. In contrast, we integrate KANs directly as feed-forward layers in a standalone SwinT encoder, making ours the first pure Transformer-based encoder with KANs for medical image segmentation.
@@
It is challenging to scale KANs in deep architectures, as vanilla spline-based KANs are computationally inefficient, memory-intensive, and difficult to train [15]. To address this, we adopt rational base functions and the Group Rational KAN (GR-KAN) design from KAT [28] to build a more robust Swin Transformer–based architecture. Our proposed model, UKAST (U-shaped segmentation architecture combining KANs with Swin Transformers), mitigates the computation and training inefficiencies of vanilla KANs, yielding a data-efficient segmentation framework with lower FLOPs and comparable parameters as to state-of-the-art (SOTA) SwinUNETR. Our UKAST achieves SOTA performance across diverse modalities of medical image segmentation benchmarks, consistently surpassing both CNN- and Transformer-based baselines. It delivers improved segmentation accuracy in two 2D and even greater improvements in two 3D datasets with approximately the same computational costs in both FLOPs and model parameters. We demonstrate that UKAST’s data efficiency outperforms that of ViTs and SwinTs across various data-scarce settings in both 2D and 3D tasks.
@@ 2.1 UKAST
Given an input 2D medical image of size $H\times W$ with $C$ number of channels, $\mathcal{X}\in\mathbb{R}^{C\times H\times W}$ (for 3D volumes, we operate on their 2D slices), we divide $\mathcal{X}$ into non-overlapping patches of size $H^{\prime}\times W^{\prime}$ each. Every patch is projected into a $C^{\prime}$-dimensional embedding space via a linear layer, forming a sequence of input tokens. The input tokens are partitioned into local windows of size $M\times M$ each in the UKAST encoder, which comprises hierarchical Swin Transformer blocks with Group-KAN feed-forward layers using rational base functions. Within each window, self-attention (W-MSA) is computed to capture local dependencies, where residual convolutions (RC) [9] are also added. To enable cross-window interactions, the subsequent layer applies shifted window attention (SW-MSA), balancing local modeling with global context aggregation. This hierarchical design progressively extracts multi-scale features, encoding both fine-grained details and broader structural information, as follows.
@@ 2.2 Rational Base Functions
We use rational functions as the base functions for KAN layers following KAT [28], rather than B-splines. Specifically, each edge function $\phi(x)$ is parameterized as a rational over polynomials $P(x),Q(x)$ of orders $m$ and $n$ respectively, using the Safe Padé Activation Unit (PAU) [17], as:
@@ 2.3 Group Rational KANs
To enhance the feed-forward pathways, we replace the conventional MLP blocks (used in SwinUNETR) with Group Rational KANs (GR-KANs) following the design of KAT [28]. Instead of using fixed activations, KAN layers learn rational function expansions, enabling expressive non-linear mappings while retaining GPU efficiency. Group-KAN reduces parameter cost by sharing base functions across neurons. Note that a key limitation of the original KAN is that it requires a unique activation function for each input-output pair, leading to a parameter count of $d_{in}\times d_{out}$ and high computational overhead. To overcome this, GR-KAN divides the input channels $d_{in}$ into $g$ groups, sharing rational function parameters within each group. This reduces the number of unique functions needed from $d_{in}\times d_{out}$ to only $g$, attaining significant savings in both parameters and FLOPs. Formally, given an input vector $\mathbf{x}$, GR-KAN applies
@@ 3 Experimental Setup
Datasets: We evaluate our new method on 4 medical image segmentation datasets spanning different modalities: polyp tissue endoscopy (Kvasir-SEG [11]), skin lesion dermoscopy (ISIC 2017 [5]), abdominal CT (BCV 2016 [12]), and cardiac MRI (MMWHS 2017 [29]). The training sets were randomly sampled and contain 700 images, 2000 images, 24 volumes (1,982 2D slices), and 16 volumes (3,834 2D slices), respectively. For data efficacy experiments, we use 10%, 25%, 50%, and 100% of the training data. The corresponding test sets include randomly selected 300 images, 600 images, 6 volumes (603 2D slices), and 4 volumes (973 2D slices).
Baselines: We evaluate three CNN-based models (U-Net [20], U-Net v2 [18], UNeXt [24]), two Transformer-based models (UNETR [8], SwinUNETR [23]), and one KAN-based model (UKAN [13]). We also construct a baseline, UKAT, by replacing the ViT [6] encoder in UNETR with the KAT encoder [28]. All models are implemented in their smallest or comparable configurations to ensure fairness. For SwinUNETR and our proposed UKAST, we additionally incorporate residual convolution blocks in the encoder stages as used in [9].
Implementation Details: The implementations utilize and MONAI [3]. The models were trained from scratch using the AdamW optimizer (initial learning rate = 0.0002, weight decay = 0.001) with a cosine annealing learning-rate schedule. The training used a combined Dice and cross-entropy loss on a single NVIDIA A10 GPU for 400 epochs with a batch size of 24. During training, $320\times 320$ random crops were extracted and augmented with horizontal and vertical flips, random $90^{\circ}$ rotations, and Gaussian noise, with masks transformed identically. At inference, predictions were generated using overlapping patches with 50% overlap.